# thanks: These three authors contributed equallythanks: These three authors contributed equallythanks: These three authors contributed equally

### v1

@@ Abstract
Antiferromagnetic resonance in a bulk $\alpha$-MnTe crystal is investigated using both frequency-domain and time-domain THz spectroscopy techniques. At low temperatures, an excitation at the photon energy of 3.5 meV is observed and identified as a magnon mode through its distinctive dependence on temperature and magnetic field. This behavior is reproduced using a simplified model for antiferromagnetic resonance in an easy-plane antiferromagnet, enabling the extraction of the out-of-plane component of the single-ion magnetic anisotropy reaching $(40\pm 10)$ $\mu$eV.
@@ I Introduction
Manganese telluride (MnTe) is a well-known antiferromagnetic semiconductor [1] with electric and magnetic properties that make it relevant for various applications, including thermoelectrics [2] or spintronics [3]. A renewed impetus to the research on MnTe – in particular, on the NiAs-type polymorph, conventionally referred to as $\alpha$-MnTe – came along with recent experimental observations [4, 5, 6, 7, 8, 9] comprising the anomalous Hall effect (AHE) [10], anisotropic magnetoresistance [11], and related phenomena at finite frequencies [7]. It is also the first collinear magnetic system where splitting of magnonic bands related to mode chirality [12] has been experimentally demonstrated [13]. This splitting is the hallmark of so-called altermagnets [14, 15, 16]: the broken parity-time ($PT$) symmetry (explained in note  [^1] Combined spatial and time inversion $T$. In addition to this symmetry, translations combined with $T$ also have to be absent among the symmetries. It should be stressed that non-collinear magnets can, of course, also have broken $PT$ symmetry.) of $\alpha$-MnTe is manifested by spin splitting of electronic bands [18, 19] that does not rely on spin-orbit coupling (SOC).
Bulk $\alpha$-MnTe displays an indirect band gap around 1.3 eV at the liquid helium temperature [20], with the conduction band minimum in the center of the Brillouin zone and valence band maxima close to the $A$ point [21]. Somewhat higher values, up to 1.5 eV, were reported for thin presumably strained epitaxial layers at room temperature [4]. Regarding magnetic structure and properties, $\alpha$-MnTe is an antiferromagnet with an ordering temperature $T_{N}$ close to 310 K [22]. The magnetic anisotropy is biaxial with a considerably stronger out-of-plane component which aligns the Mn2+ spins in the hexagonal planes [23, 24]. Within each plane, the spins are ordered ferromagnetically. The coupling between adjacent planes is antiferromagnetic (Fig. 1a).
@@
The dispersion of magnon excitations in $\alpha$-MnTe was mapped by means of inelastic neutron scattering experiments and modeled, using linear spin wave (LSW) theory, by Szuszkiewicz et al. [25]. In their model, they considered an easy-plane type of the antiferromagnetic order and therefore, as is known in NiF2 [26], the double degeneracy of magnon dispersion is lifted at the center of the magnetic Brillouin zone ($k=0$ magnons). However, experimentally, only a single magnon mode at the energy of $2.6\pm 0.8$ meV ($0.65\pm 0.20$ THz) was resolved at the $\Gamma$ point [25].
In this study, we present antiferromagnetic resonance (AFMR) experiments in the THz spectral range conducted on bulk $\alpha$MnTe across a wide range of temperatures and applied magnetic fields. The observed behavior is consistent with theoretical predictions for an easy-plane (hard-axis) antiferromagnet, based on the LSW theory. Our modeling allows us to estimate the out-of-plane anisotropy – the parameter that defines the magnetic properties of MnTe to a great extent. This refines the value reported in the previous study using inelastic neutron scattering [25].
@@ II Theoretical model
The antiferromagnetic resonance is a phenomenon widely explored in solid-state physics, approached theoretically both at the classical and quantum levels [27, 28, 29]. Here we summarize theoretical expectations for the AFMR in an easy-plane antiferromagnet, which can be taken as the first approximation for $\alpha$-MnTe. Hence, we neglect the in-plane anisotropy as well as the (higher-order) Dzyaloshinskii-Moriya interaction which are both relevant for MnTe [6], but they imply energy scales at least one order of magnitude smaller than the out-of-plane anisotropy. Similarly, the proposed model, see Appendix A for details, disregards magnon-magnon interactions. Their inclusion should not lead to a qualitative change in the dispersion of magnons, apart from its broadening [30]. Hence, our model should be sufficient for a description of the dispersion profile.
@@
Clearly, the above Hamiltonian implies three distinct energy scales: the Zeeman energy $E_{Z}=g\mu_{B}B$, the magnetic anisotropy energy $E_{\mathrm{an}}=DS$ (due to both the SOC affecting the band structure and the dipole-dipole interaction [31]), and the effective exchange energy $E_{J}=\sum_{n}J_{n}{\cal Z}_{n}S$, where $n$ runs over the neighboring inter-sublattice spins ($n=1$ and 3), with the exchange coupling strength of $J_{n}$ and the corresponding coordination number ${\cal Z}_{n}$ (${\cal Z}_{1}=2$, ${\cal Z}_{3}=12$).
@@
The magnon dispersion, including the $k=0$ modes relevant for our optical and magneto-optical experiments, is often obtained using the LSW theory [25, 29]. In antiferromagnets with easy-plane magnetic anisotropy and two spins per unit cell, such as MnTe (see Fig. 1a), we expect the degeneracy of the two magnon modes to be lifted throughout the Brillouin zone. At the $\Gamma$ point, the separation between them scales with the strength of the magnetic anisotropy. At non-zero momenta, interestingly, a part of the splitting may be associated with the altermagnetic nature of MnTe [12, 13]. In the following, we will refer to the lower and upper magnon branches at $k=0$ as the $\beta$ and $\alpha$ modes, respectively.
The LSW theory predicts the energies of both magnon modes, $\hbar\omega_{\alpha}(B)$ and $\hbar\omega_{\beta}(B)$, for any strength and direction of the applied magnetic field. Particularly simple expressions are obtained in two limiting cases: for $\mathbf{B}\perp c$ ($\theta_{B}=\pi/2$), when $B$ exceeds the corresponding spin-flop fields ($B_{\mathrm{sf}}$ below 4 T, see Refs. [24, 32]) and for ${\mathbf{B}}\,\|\,c$ ($\theta_{B}=0$). In the following, we will call these cases “spin-flop” (SF) and “even canting” (EC) configurations, respectively.
In our simplified Hamiltonian for spins in MnTe (1), the $k=0$ magnon energies depend on the applied magnetic field ($E_{Z}\propto B$) as
@@
$\hbar\omega^{\mathrm{EC}}_{\alpha}$ $=\sqrt{4E_{\mathrm{an}}E_{J}+\frac{E_{J}E_{Z}^{2}(E_{J}-E_{\mathrm{an}})}{(E_{J}+E_{\mathrm{an}})^{2}}},$ (3)
@@
At $B=0$, these results agree with Eqs. 23 and 24 of Ref. [29] in the limit of a vanishing in-plane anisotropy. At a non-zero magnetic field, but still below $B_{\mathrm{sf}}$, the energies of magnon modes strongly depend on the initial angle between magnetic moments and the applied (in-plane) magnetic field. This case was treated theoretically [33] and studied experimentally in easy-plane antiferromagnets, e.g., in the recent magneto-Raman scattering experiment on NiPS3 [34], but it is not relevant for experimental data presented in this work.
In $\alpha$-MnTe, one may expect $E_{J}\gg E_{\mathrm{an}}$ [25, 13] which allows us to plot, in a qualitative way, the expected $B$ dependence of $k=0$ modes, see Fig. 1b,c. In the SF and EC configurations, we find distinctively different behavior. In the EC case, the $\alpha$ mode undergoes a profound blueshift, approximately quadratic at low magnetic fields, while the $\beta$ mode stays at zero, unless effects due to the considerably weaker in-plane anisotropy are considered, see e.g., Ref. [29]. In the SF case, the $\alpha$ mode undergoes a weak redshift with $B$, while the $\beta$ mode approaches the spin resonance of a free electron. Unfortunately, as discussed in the following, only the $\alpha$ mode is observed in our THz experiments.
Figure 3: (a) Temperature-differential reflectivity of $\alpha$-MnTe (sample B) as a function of frequency and temperature measured using the time-domain technique, compared with results of the theoretical model (b) explained in the main text and Appendix E. (c) Experimentally determined magnon energy as a function of temperature for samples B and A. Observed decrease in energy is compared with universality class models implying $(T/T_{N})^{\zeta}$ scaling [35]. (d) Electrical resistivity of sample B as a function of $T$ with the theoretical fit described in Appendix E.
@@ III Experimental
Bulk MnTe samples, studied in this work, were prepared in two batches using different growth methods. Sample A was grown using the self-flux technique. Pure manganese (99.9998%) and tellurium (99.9999%) in the molar composition Mn33Te67 were placed in an alumina (99.95%) crucible and, together with a catch crucible filled with quartz wool, sealed in a fused-silica tube under vacuum. The sample was heated up to 1050 C and then cooled down to 760 C for four days. At 760 C, the sample was quickly placed in a centrifuge, separating the crystals from the remaining melt. The crystals were flat plates with lateral dimensions of several millimeters and with the thickness of several hundred of microns. Bulk MnTe crystals grown under identical conditions had been characterized using powder and single-crystal X-ray techniques, magnetic susceptibility measurements, and XPS spectroscopy in Ref. [18].
@@
Four different methods of THz spectroscopy were employed to gather experimental data presented in this paper: (i) Fourier-transform magneto-spectroscopy at low temperatures; (ii) the time-domain technique employed at $B=0$ in a broad range of temperatures with a complementary set of data collected under magnetic field applied at low temperatures; (iii) magnetic-circular-dichroism (MCD) technique at selected THz frequencies and varying temperatures; and (iv) high-frequency electron-spin resonance (ESR). These experimental techniques are described in detail in Appendix D.
Figure 4: A color map of magneto-transmission taken on sample A at liquid helium temperature using (a) superconducting and (b) resistive coils, with the magnetic field applied along the $c$ axis. The magneto-transmission $T_{B}$ is normalized by the transmission averaged over the whole field range. The extracted positions of resonances with a step of 2 T are plotted as blue circles in (c), complemented with data obtained by the time-domain THz and ESR techniques (green and dark orange circles, respectively), see Appendix F for the latter. The size of the circles corresponds to the error bar. For the dashed line, see the main text. Panel (d) displays the magneto-transmission of sample A, probed by the THz/FIR laser. The magnetic-field sweep at a temperature of $T=200$ K was obtained using circularly polarized radiation with a photon energy of $\hbar\omega=2.87$ meV. The temperature sweeps, shown in the inset, were measured at several magnetic fields using a linearly polarized laser beam with a photon energy of 2.71 meV. Panel (e) shows low-temperature magneto-transmission plotted as a color map, for $B$ applied at the angle of $\theta_{B}=50^{\circ}$ with respect to the $c$ axis.
@@ IV Results and discussion
Let us start the discussion of our experimental results with the transmission spectrum of sample A measured using the Fourier-transform technique at $B=0$ and $T=4.2$ K, see the blue line in Fig. 2. In the THz range, the transmission gradually decreases with photon energy and vanishes above 9 meV due to the onset of the reststrahlen band. In the explored spectral window, the transmission spectrum is weakly modulated by an interference pattern. We observe only a single absorption mode marked by the vertical arrow in Fig. 2. The same excitation is also visible in the extinction coefficient $\kappa$ extracted using the time-domain THz spectroscopy, see the orange line in Fig. 2. As justified a posteriori, by the characteristic temperature and magnetic-field dependence, this is the AFMR signal of MnTe. Equivalently, we may call it the $k=0$ magnon mode that corresponds to the upper dispersion branch ($\alpha$ mode). At low temperatures, the mode has an energy of $E_{\alpha}=\hbar\omega_{\alpha}^{B=0}=(3.5\pm 0.1)$ meV. For comparison, the magnon energy reported for an epitaxial layer of bulk cubic MnTe is larger, reaching nearly 4.3 meV at low temperatures [36, 37].
In Fig. 3a, we present, in the form of a color plot, $T$-differential reflection spectra of sample B, measured using THz time-domain spectroscopy (temperature-differential approach is explained in Appendix D). The characteristic redshift of the magnon mode with $T$ is directly visible in the raw data and we performed a more detailed analysis to extract its spectral position. Apart from the magnon excitation, the studied optical response also depends on free carriers (holes), the dynamics of which display a marked temperature dependence on both the density and scattering rate. The latter is manifested by a pronounced minimum of electrical resistivity (Fig. 3d and Appendix B) at $T\approx 50$ K that coincides with the minimum of optical transmission in the THz spectral range. This minimum separates the low- and high-temperature regimes, where the temperature-driven changes in resistivity are mainly due to the freeze-out of free carriers on shallow acceptor centers and due to thermally activated scattering, respectively. We have modeled this behavior using a simple Drude-type approach which provided us with parameters required for a reliable fit of our THz reflectivity data, see Appendix E.
@@
The extracted positions of the magnon resonance are plotted in Fig. 3b, complemented by several control points obtained on sample A using an independent time-domain THz setup. With increasing $T$, the observed mode exhibits a pronounced redshift, down to 1.4 meV at room temperature. Such a redshift is characteristic of magnon excitations in magnetic materials when approaching, from below, the ordering temperature, see, e.g., Refs. [38, 39]. In Fig. 3b, we modeled this behavior using the formula $E_{\alpha}(T)=E_{\alpha}[1-c(T/T_{N})^{\zeta}]$, with $\zeta$ and $c$ as the fitting parameters. Consistent with Refs. [40, 35], this yielded values of $\zeta=(2.5\pm 0.1)$ and $c=(0.68\pm 0.05)$.
@@
Our optical measurements on $\alpha$-MnTe in an external magnetic field are summarized in Fig. 4. The basic set of the collected data, with $\mathbf{B}$ applied along the $c$ axis of MnTe (EC configuration) is presented in panel (a). The observed magnetic excitation, exhibiting a pronounced blueshift, approximately quadratic in $B$, can be directly associated with the $\alpha$ mode. Despite a lower signal-to-noise ratio, the monotonic blueshift is well visible also in the high-field part of the data, see panel (b), and continues up to the highest magnetic field applied (33 T). The extracted energies of the $\alpha$ mode are plotted in panel (c) together with the low-field points obtained using the time-domain THz spectroscopy and ESR techniques.
@@
Interestingly, another field-dependent spectral feature emerges at lower photon energies, see Figs. 4a and b and orange circles in Fig. 4c. One could invoke the possibility that this extra line is the lower magnon branch ($\beta$ mode). In fact, also the $\beta$ mode may acquire a non-zero energy and gain a non-trivial dependence on $B$ when the in-plane component of anisotropy is not fully negligible. However, a closer inspection of our data collected on other MnTe samples from batch A excludes this scenario. In contrast to the $\alpha$ line, which is consistently present in all our measurements, the lower energy line is not manifested in all samples. Presumably, this reflects the varying hole density among the samples studied. In addition, the lower-energy line broadens significantly with $B$. This is considerably different from behaviour of the $\alpha$ line and represents another argument against the magnon origin of this extra line.
@@
To find an interpretation for this extra line, let us recall the nearly linear dependence of its position on the applied magnetic field, see Fig. 4c. This may remind us of conventional cyclotron resonance (CR) of free charge carriers [41]. However, at the same time, the zero-field extrapolation yields a nonzero energy. Such behavior was observed in previous CR studies of weakly bound electrons in bulk semiconductors and their heterostructures, see, e.g., Refs. [42, 43, 44, 45]. In fact, the presence of electrons localized at low temperatures is consistent with a sharp drop in electrical resistance observed with increasing $T$, see Fig. 5. The slope of the extra line would imply a rather high effective mass, $m_{\mathrm{eff}}\approx m_{0}$. This is, however, greater than the previously reported experimental values ($\approx 0.25m_{0}$ confirmed also by ab initio calculations, see Sec. 2 of the Supplementary information in Ref. [21]) based on zero-field optical studies at liquid nitrogen temperature [46, 47].
Let us now turn back to the $\alpha$ mode (Figs. 4a-c) and examine its other experimentally probed characteristics, the partial circular polarization at $B>0$, see Fig. 4d. It was observed using the THz MCD technique, which operates at a fixed laser frequency with a well-defined circular polarization, and follows the intensity of the transmitted radiation as a function of $B$. The partial circular polarization of the $\alpha$ mode is manifested by a profound difference in the observed absorption when the light propagates parallel or antiparallel to the direction of the magnetic field.
Magnetic circular dichroism in MnTe is a complex phenomenon, sensitive to particular symmetries present (or broken), see, e.g. Ref. [7], nevertheless, the partial circular polarization of the $\alpha$ mode at $B>0$ can be seen just as an effect of the $B$-induced magnetization. In the ${\mathbf{B}}\|c$ configuration, this appears due to spin canting away from the hexagonal plane (the EC regime). The angle between the spin and hexagonal planes reaches a few degrees in the highest applied $B$, $\arcsin[E_{Z}/(2E_{J})]$ (6), for the expected strength of the exchange coupling [25]. Note that the experiment with circularly polarized radiation was performed at higher temperatures ($T=200$ K) at which it was possible to achieve resonance between the field-dependent $\alpha$ mode and one of the (discrete) emission lines of the gas laser. This temperature-driven tuning of the magnon energy, at several values of the magnetic field, is demonstrated in the inset of Fig. 4d.

### v2

@@ Abstract
Antiferromagnetic resonance in a bulk $\alpha$-MnTe crystal is investigated using both frequency-domain and time-domain THz spectroscopy techniques. At low temperatures, an excitation at the photon energy of $(3.5\pm 0.1)$ meV is observed and identified as a magnon mode through its distinctive dependence on temperature and magnetic field. This behavior is reproduced using a simplified model for antiferromagnetic resonance in an easy-plane antiferromagnet. The results of our experiments, when compared to exchange constants established in the literature, allow us to extract the out-of-plane component of the single-ion magnetic anisotropy reaching $(40\pm 10)$ $\mu$eV.
@@ I Introduction
Manganese telluride (MnTe) is a well-known antiferromagnetic semiconductor [1] with electric and magnetic properties that make it relevant for various applications, including thermoelectrics [2] or spintronics [3]. A renewed impetus to the research on MnTe – in particular, on the NiAs-type polymorph, conventionally referred to as $\alpha$-MnTe – came along with recent experimental observations [4, 5, 6, 7, 8, 9] comprising the anomalous Hall effect (AHE) [10], anisotropic magnetoresistance [11], and the magnetic circular dischroism (MCD) effect at optical frequencies [7]. It is also the first collinear magnetic system where splitting of magnonic bands related to mode chirality [12] has been experimentally demonstrated [13]. This splitting is the hallmark of so-called altermagnets [14, 15, 16]. The broken parity-time ($PT$) symmetry, characteristic of altermagnets (explained in note  [^1] Combined spatial and time inversion $T$. In addition to this symmetry, translations combined with $T$ also have to be absent among the symmetries. It should be stressed that non-collinear magnets can, of course, also have broken $PT$ symmetry.), is manifested by spin splitting of electronic bands [18, 19] that does not rely on spin-orbit coupling (SOC).
Bulk $\alpha$-MnTe displays an indirect band gap around 1.3 eV at the liquid helium temperature [20], with the conduction band minimum in the center of the Brillouin zone and valence band maxima close to the $A$ point [21]. Somewhat higher values, up to 1.5 eV, were reported for thin presumably strained epitaxial layers at room temperature [4]. Regarding magnetic structure and properties, $\alpha$-MnTe is an antiferromagnet with an ordering temperature $T_{N}$ close to 310 K [22]. The magnetic anisotropy is biaxial with a considerably stronger out-of-plane component which aligns the Mn2+ spins in the hexagonal planes [23, 5, 24]. Within each plane, the spins are ordered ferromagnetically. The coupling between adjacent planes is antiferromagnetic (Fig. 1a).
@@
The dispersion of magnon excitations in $\alpha$-MnTe was mapped by means of inelastic neutron scattering experiments and modeled, using linear spin wave (LSW) theory, by Szuszkiewicz et al. [25]. In their model, they considered an easy-plane type of the antiferromagnetic order and therefore, as is known in NiF2 [26], the double degeneracy of magnon dispersion is lifted at the center of the magnetic Brillouin zone ($k=0$ magnons). Experimentally, a single magnon mode at the energy of $(2.6\pm 0.8)$ meV, most likely the upper magnon branch, was reported at the $\Gamma$ point in [25]. In a more recent neutron scattering experiment by Liu et al. [13], a bit higher energy of this mode was concluded, close to 4 meV, based on data modeling in the framework of the LSW theory.
The energies and splitting of long-wavelength magnon modes are largely determined by the orientation and strength of single-ion magnetic anisotropy which plays a key role in defining the type and stability of ordered phases, not only in antiferromagnetic materials (see, e.g., [27, 28]). Valuable insights in magnetic anisotropy can thus be obtained by experimentally probing the $k=0$ magnons. For this purpose, magnetic resonance is the method of choice, as it allows for precise determination of magnon gaps, typically with greater accuracy than inelastic neutron scattering.
In this study, we present antiferromagnetic resonance (AFMR) experiments in the THz spectral range conducted on bulk $\alpha$-MnTe across a wide range of temperatures and applied magnetic fields. The observed behavior is consistent with theoretical predictions for an easy-plane (hard-axis) antiferromagnet, based on the LSW theory. Our modeling allows us to estimate the out-of-plane anisotropy and compare it with the results of previous studies using inelastic neutron scattering [25, 13].
@@ II Theoretical model
The AFMR is a phenomenon widely explored in solid-state physics, approached theoretically at the classical and quantum levels [29, 30, 28]. It refers to resonant absorption of light in antiferromagnets that emerges due to the precession of exchange-coupled spin sub-lattices subjected to an externally applied magnetic field. Alternatively, the AFMR mode can be viewed as an optically driven excitation of the $k=0$ magnon modes.
Here we summarize theoretical expectations for the AFMR in an easy-plane antiferromagnet, which can be taken as the first approximation for $\alpha$-MnTe. Hence, we neglect the in-plane anisotropy as well as the (higher-order) Dzyaloshinskii-Moriya interaction which are both relevant for $\alpha$-MnTe [6], but they imply energy scales at least one order of magnitude smaller than the out-of-plane anisotropy. Similarly, the proposed model, see Appendix A for details, disregards magnon-magnon interactions. Their inclusion should not lead to a qualitative change in the dispersion of magnons, apart from its broadening [31]. Hence, our model should be sufficient for a description of the dispersion profile.
@@
The above Hamiltonian implies three distinct energy scales: the Zeeman energy $E_{Z}=g\mu_{B}B$, the magnetic anisotropy energy $E_{\mathrm{an}}=DS$ (due to both the SOC affecting the band structure and the dipole-dipole interaction [32]), and the effective exchange energy $E_{J}=\sum_{n}J_{n}{\cal Z}_{n}S$, where $n$ runs over the neighboring inter-sublattice spins ($n=1$ and 3), with the exchange coupling strength of $J_{n}$ and the corresponding coordination number ${\cal Z}_{n}$ (${\cal Z}_{1}=2$, ${\cal Z}_{3}=12$). Let us note that the intra-sublattice coupling $J_{2}$ does not enter $E_{J}$ in the macrospin representation, used to calculate the ground state, see Appendix A.
@@
The magnon dispersion, including the $k=0$ modes relevant for our optical and magneto-optical experiments, can be obtained using the LSW theory [25, 28]. In antiferromagnets with easy-plane magnetic anisotropy and two spins per unit cell, such as $\alpha$-MnTe (see Fig. 1a), we expect the degeneracy of the two magnon modes to be lifted throughout the Brillouin zone. At the $\Gamma$ point, the separation between them scales with the strength of the magnetic anisotropy. At non-zero momenta, interestingly, a part of the splitting may be associated with the altermagnetic nature of $\alpha$-MnTe [12, 13]. In the following, we will refer to the lower and upper magnon branches at $k=0$ as the $\beta$ and $\alpha$ modes, respectively.
The LSW theory predicts the energies of both magnon modes, $\hbar\omega_{\alpha}(B)$ and $\hbar\omega_{\beta}(B)$, for any strength and direction of the applied magnetic field. Particularly simple analytic expressions are obtained in two limiting cases: for $\mathbf{B}\perp c$ ($\theta_{B}=\pi/2$), when $B$ exceeds the corresponding spin-flop fields ($B_{\mathrm{sf}}$ below 4 T, see Refs. [5, 33]) and for ${\mathbf{B}}\,\|\,c$ ($\theta_{B}=0$). In the following, we will call these cases “spin-flop” (SF) and “even canting” (EC) configurations, respectively.
In our simplified Hamiltonian for spins in $\alpha$-MnTe (1), the $k=0$ magnon energies depend on the applied magnetic field ($E_{Z}\propto B$) as
@@
$\hbar\omega^{\mathrm{EC}}_{\alpha}$ $=\sqrt{4E_{\mathrm{an}}E_{J}+\frac{E_{Z}^{2}(1-E_{\mathrm{an}}/E_{J})}{(1+E_{\mathrm{an}}/E_{J})^{2}}},$ (3)
@@
The case of the SF configuration, but at magnetic fields below $B_{\mathrm{sf}}$, was treated theoretically, e.g., in Ref.  [34]. The energies of both magnon modes are expected to depend strongly on the initial angle between magnetic moments and the applied (in-plane) magnetic field. We did not observe such effects in $\alpha$-MnTe, but they were studied in some other easy-plane systems with higher spin-flop fields, e.g., in NiPS3 [35]. Let us also note that, at $B=0$, the above results agree with Eqs. 23 and 24 of Ref. [28] in the limit of a vanishing in-plane anisotropy. The nature of $\alpha$ and $\beta$ modes can be visualized in terms of a classical precession of sub-lattice magnetizations (see Fig. 5 in [28]). In the aforementioned limit, the corresponding trajectories become highly elliptical and linear, respectively.
In Figs. 1b and c, we plot the expected $B$ dependence of the $k=0$ magnon modes. Both energies are normalized by the energy of the $\alpha$ mode at $B=0$ ($E_{\alpha}$). The realistic ratio $E_{\mathrm{an}}/E_{J}=0.003$, justified a posteriori in this work, was used. We find distinctively different behaviour in the SF and EC configurations. In the EC case, the $\alpha$ mode undergoes a profound blueshift, approximately quadratic at low magnetic fields, while the $\beta$ mode stays at zero, unless effects due to the considerably weaker in-plane anisotropy are considered, see e.g., Ref. [28]. In the SF case, the $\alpha$ mode undergoes a weak redshift with $B$, while the $\beta$ mode approaches the spin resonance of a free electron. Unfortunately, as discussed in the following, only the $\alpha$ mode is observed in our THz experiments.
Figure 3: (a) Temperature-differential reflectivity of $\alpha$-MnTe (sample B) as a function of frequency and temperature measured using the time-domain technique, compared with results of the theoretical model (b) explained in the main text and Appendix D. (c) Experimentally determined magnon energy as a function of temperature for samples B and A. Observed decrease in energy is compared with universality class models implying $(T/T_{N})^{\zeta}$ scaling [36]. (d) Electrical resistivity of sample B as a function of $T$ with the theoretical fit described in Appendix D.
@@ III Experimental
Bulk $\alpha$-MnTe samples, studied in this work, were prepared in two batches using different growth methods. Sample A was grown using the self-flux technique. Pure manganese (99.9998%) and tellurium (99.9999%) in the molar composition Mn33Te67 were placed in an alumina (99.95%) crucible and, together with a catch crucible filled with quartz wool, sealed in a fused-silica tube under vacuum. The sample was heated up to 1050 C and then cooled down to 760 C for four days. At 760 C, the sample was quickly placed in a centrifuge, separating the crystals from the remaining melt. The crystals were flat plates with lateral dimensions of several millimeters and a thickness of several hundred microns. Bulk $\alpha$-MnTe crystals grown under identical conditions had been characterized using powder and single-crystal X-ray techniques, magnetic susceptibility measurements, and XPS spectroscopy in Ref. [18].
@@
Four different methods of THz spectroscopy were employed to gather experimental data presented in this paper: (i) Fourier-transform magneto-spectroscopy at low temperatures (Figs. 2, 4 and 5); (ii) the time-domain technique employed at $B=0$ (Figs. 2 and 3) in a broad range of temperatures with a complementary set of data collected under magnetic field applied at low temperatures (Fig. 4c); (iii) magnetic-circular-dichroism (MCD) technique at selected THz frequencies and varying temperatures (Fig. 6); and (iv) high-frequency electron-spin resonance (ESR) in Figs. 4c and 8. These experimental techniques are described in detail in Appendix C.
Figure 4: A color map of magneto-transmission taken on sample A at liquid helium temperature using (a) superconducting and (b) resistive coils, with the magnetic field applied along the $c$ axis. The magneto-transmission $T_{B}$ is normalized by the transmission averaged over the whole field range. The extracted positions of resonances with a step of 2 T are plotted as blue circles in (c), complemented with data obtained by the time-domain THz and ESR techniques (green and dark orange circles, respectively), see Appendix E for the latter. The size of the circles corresponds to the error bar. For the dashed line, see the main text.
@@ IV Results and discussion
Let us start the discussion of our experimental results with the transmission spectrum of sample A measured using the Fourier-transform technique at $B=0$ and $T=4.2$ K, see the blue line in Fig. 2. In the THz range, the transmission gradually decreases with photon energy and vanishes above 9 meV due to the onset of strong phonon absorption. In the explored spectral window, the transmission spectrum is weakly modulated by an interference pattern. We observe only a single absorption mode marked by the vertical arrow in Fig. 2. The same excitation is also visible in the extinction coefficient $\kappa$ extracted using the time-domain THz spectroscopy (at $T=7$ K), see the orange line in Fig. 2. As justified a posteriori, by the characteristic temperature and magnetic-field dependence, this is the AFMR signal of $\alpha$-MnTe. Equivalently, we may call it the $k=0$ magnon mode that corresponds to the upper dispersion branch ($\alpha$ mode). At low temperatures, the mode has an energy of $E_{\alpha}=\hbar\omega_{\alpha}^{B=0}=(3.5\pm 0.1)$ meV. For comparison, the magnon energy reported for an epitaxial layer of bulk cubic MnTe is larger, reaching nearly 4.3 meV at low temperatures [37, 38].
In Fig. 3a, we present, in the form of a color plot, temperature-differential reflection spectra of sample B, measured using THz time-domain spectroscopy (see Appendix C) in the range of $T=10-300$ K. The reflectivity configuration was used, as the sample opacity in the THz range increases significantly with temperature. This is clearly manifested in the optical data by a profound change in the interference pattern around 30 K (Fig. 3a). We interpret this as due to free charge carriers (holes), excited thermally from acceptor states. This is in line with the rapid decrease in dc electrical resistivity observed when the temperature increases from 10 to 50 K (Fig. 3d).
Even though the $\alpha$ magnon mode and its characteristic redshift with $T$ is directly traceable in the raw data (Fig. 3a), we performed a more detailed analysis to extract its position, width, and strength. To describe the contribution of holes to THz conductivity, thus reducing the number of free parameters in the modeling of our time-domain reflectivity data, we reproduced the temperature dependence of dc resistivity by an empirical function, see Appendix D for details.
@@
The extracted positions of the magnon resonance are plotted in Fig. 3c, complemented by several control points obtained on sample A using an independent time-domain THz setup. With increasing $T$, the observed mode exhibits a pronounced redshift, down to 1.4 meV at room temperature. Such a redshift is characteristic of magnon excitations in magnetic materials when approaching, from below, the ordering temperature, see, e.g., Refs. [39, 40]. In Fig. 3c, we modeled this behavior using the formula $E_{\alpha}(T)=E_{\alpha}[1-c(T/T_{N})^{\zeta}]$, with $\zeta$ and $c$ as the fitting parameters. Consistent with Refs. [41, 36], this yielded values of $\zeta=(2.5\pm 0.1)$ and $c=(0.68\pm 0.05)$ for $T_{N}=310$ K. Our modeling procedure also provided us with the width and oscillator strength (magnetic susceptibility) of the $\alpha$ mode. At low temperatures, these reach $\gamma_{m}=(20\pm 10)$ $\mu$eV and $\chi=(6\pm 2)\times 10^{-4}$, respectively. The details of modeling and its results are discussed in Appendix D, including the temperature dependence.
@@
Our optical measurements on $\alpha$-MnTe in an external magnetic field are summarized in Fig. 4. The basic set of the collected data, with $\mathbf{B}$ applied along the $c$ axis of $\alpha$-MnTe (EC configuration) is presented in panel (a). Magnetic excitation, identified above as the $\alpha$ mode, undergoes a pronounced blueshift that is approximately quadratic in $B$. Despite a lower signal-to-noise ratio, the monotonic blueshift is well visible also in the high-field part of the data, see panel (b), and continues up to the highest magnetic field applied (33 T). The extracted energies of the $\alpha$ mode are plotted in panel (c) together with the low-field points obtained using the time-domain THz spectroscopy and ESR techniques.
@@
Interestingly, another field-dependent spectral feature emerges at lower photon energies, see Figs. 4a and b and orange circles in Fig. 4c. One could invoke the possibility that this extra line is the lower magnon branch ($\beta$ mode). In fact, also the $\beta$ mode may acquire a non-zero energy when the in-plane component of anisotropy is not fully negligible. However, a closer inspection of our data collected on other $\alpha$-MnTe samples from batch A excludes this scenario. In contrast to the $\alpha$ line, which is consistently present in all our measurements, the lower energy line is not manifested in all samples. Presumably, this reflects the varying hole density among the samples studied. In addition, the lower-energy line broadens significantly with $B$. This is considerably different from behaviour of the $\alpha$ line and represents another argument against the magnon origin of this extra line.
@@
To find an interpretation for this extra line, let us recall the nearly linear dependence of its position on the applied magnetic field, see Fig. 4c. This may remind us of conventional cyclotron resonance (CR) of free charge carriers [42]. However, at the same time, the zero-field extrapolation yields a nonzero energy. Such behavior was observed in previous CR studies of weakly bound electrons in bulk semiconductors and their heterostructures, see, e.g., Refs. [43, 44, 45, 46]. In fact, the presence of electrons localized at low temperatures is consistent with a sharp drop in electrical resistance observed with increasing $T$, see Fig. 7. The slope of the extra line would imply a rather high effective mass, $m_{\mathrm{eff}}\approx m_{0}$. This is, however, greater than the previously reported experimental values ($\approx 0.25m_{0}$) extracted from zero-field optical studies at liquid nitrogen temperature [47, 48] as well as the ab-initio estimate [21].